Thanks in advance for your assistance!
Solve for A and B in the following problems:
1)4A+B = 16, 2A = 3B =24
2) A+B=6, 2A+B=8

Answer by gsmani_iyer(201)

Check the signs in the first problem and resubmit.
2) A + B = 6 ......(1)
2A + B = 8 ......(2)
(2) - (1), we get
A = 2
By substituting the value of A in Eq.(1),
2 + B = 6
So B = 6-2 = 4
A = 2; B = 4. Ans.

In the first step, we have eliminated one variable (B) by subtracting.
In the second step, we have substituted the value of A and solved.